OpenStudy (anonymous):

An integrating factor of the ODE (e^x*sec(y)-48*tan(y))dx+dy=0 is (e^(-a*x))*cosy. Find a.

OpenStudy (anonymous):

What do you know about integrating factors?

OpenStudy (anonymous):

In the question \[e^{-ax}\cos(y)\] is the integrating factor of the ODE

OpenStudy (anonymous):

Yes, what properties does the integrating factor have?

OpenStudy (anonymous):

Integrating factor can make ODE (Ordinary Differential Equations) as an Exact differential form like as M(x,y)dx + N(x,y)dy = 0 and partial derivatives of M(x,y)|y = N(x,y)|x

OpenStudy (anonymous):

I tried to find integrating factor of above ODE using this method : \[(M_{y}-N_{x})/M =\frac{ e^xcsc(y)\tan(y)-48\sec^2(y) }{ e^xsec(y)-48\tan(y) }\] However I can't make \[e^{-ax}\cos(y)\] form.

OpenStudy (psymon):

Im not able to get it to come out either, but two corrections. First, should be e^xsec(y)tan(y) in the numerator. Second, you have it reversed for the formula. Its just \[\frac{ N _{x}-M _{y} }{ M }\] Other than that, Im pretty much in this chapter in my class, so not sure I would know much more than you x_x